I need help with part A and part B. I tried those 3 answers and none were correct.
media%2Ff7d%2Ff7da789f-51dd-4ffb-8e96-fb
HOT >パ며 tlectric Charge Content PHY × 며 Course Modules: PHY 222-E mode-myAnswers&assignment .: myemich-myorích C Secure https/ Submitted Answers ANSWER 1 ANSWER 2 ANSWER 3: F 4.7 N ANSWER 4 Pot 4.7N
Answer #1

(A) It is given that (q ) = +2.6 uC
d = 31 cm = 0.31 m
Force on charge 2 due to charge 1
kq2qi
On putting the values we get
910 2 2.6 102.610 F21 0.312
On solving we get
F21 = 1.266 N
Now force on charge 2 due to charge 3
kq2q3 2312
On putting the values we get
9*109 * 2 * 2.6 * 10-6 * 3 * 2.6 * 10-6 0.312 F23
On solving , we get
F23 = 3.798 N
Now force on charge 2 due to charge 4
kq2q4 (dv2)2 24
On putting the values we get
9 10 2 2.6 1042.6 10 F24 (0.31V2)2

On solving , we get
F24 = 2.532 N
Now all forces are shown in the figure
In the x direction
FX = -F23 -F24Cos45 = -3.798 - 2.532Cos45 = - 5.588 N
In the Y direction
FY = F24Sin45 - F21 = 2.532Sin45 - 1.266 = 0.524 N
Therefore the magnitude of the force would be
3 2
F = 5.613 N
Now the direction would be
Fy
0.524 tano_5,588
-5.357 in clockwise direction from q3 - q2 direction
Now in counterclockwise direction it would be
phi = 180 + heta
phi= 174.64 o in counter clockwise direction from q3 - q2 direction
24 F, Sin45 24 Sin45 g-2q 43-3 q 45 23 F Cos45 21 45 44-4q